OpenStudy (anonymous):

one of two numbers is two-thirds the other number. the sum of the numbers is 45. find the numbers

OpenStudy (anonymous):

x + 2/3x = 45 5/3x = 45 x = 45 *3/5 x = 27 27 and 27*2/3 = 18